OpenStudy (anonymous):

Using the given equations, find the point of intersection. If there is no point of intersection, type no points . 3 x + 2 y = 5 5 x + 2 y = 7 The lines intersect at ___________ Using the given equations, find the point of intersection. If there is no point of intersection, type no points . 3 x + 2 y = 5 5 x + 2 y = 7 The lines intersect at ___________ @Mathematics

OpenStudy (anonymous):

-3x - 2y = -5 5x + 2y = 7 ______________ -2x/2 = 2/-2 x= -1